# Khoroshkin–Tolstoy approach for quantum superalgebras

### v2

@@ Abstract.
The central object of the quantum algebraic approach to the study of quantum integrable models is the universal $R$-matrix, which is an element of a completed tensor product of two copies of quantum algebra. Various integrability objects are constructed by choosing representations for the factors of this tensor product. There are two approaches to constructing explicit expressions for the universal $R$-matrix. One is based on the quantum double construction, and the other is based on the concept of the $q$-commutator. In the case of a quantum superalgebra, we cannot use the first approach, since we do not know an explicit expression for the Lusztig automorphisms. One can use the second approach, but it requires some modifications related to the presence of isotropic roots. In this article, we provide the necessary modification of the method and use it to find an $R$-operator for quantum integrable systems related to the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$.
@@ 1. Introduction
The study of quantum integrable models is based on the constructing the corresponding integrability objects and solving the functional relations they satisfy. [^1] For the terminology used for integrability objects, we refer to the papers [1, 2, 3]. Here, the most productive method is the quantum algebraic approach. [^2] Previously, what we call quantum algebra was usually called quantum group. In fact, this object is an associative algebra, which in a sense is a deformation of the universal enveloping algebra of a Lie algebra. Nowadays, the term quantum algebra is more commonly used, and we adhere to this terminology. The central object of this approach is the universal $R$-matrix, which is an element of the completed tensor product of two copies of the quantum algebra. Various integrability objects are constructed by choosing representations for the factors of this tensor product. The consistent application of the quantum algebraic approach was initiated by Bazhanov, Lukyanov and Zamolodchikov [4, 5, 6]. They considered the quantum version of KdV theory. Subsequently, this method turned out to be efficient for studying other quantum integrable models.
@@
The general notion of a quantum algebra $\mathrm{U}_{q}(\mathfrak{g})$ was proposed by Drinfeld and Jimbo [7, 8] for the case when $\mathfrak{g}$ is a Kac–Moody algebra with a symmetrizable generalized Cartan matrix. In the report [9], Drinfeld described the quantum double construction of the universal $R$-matrix and presented an explicit expression for the quantum group $\mathrm{U}_{q}({\mathfrak{sl}_{2}})$. Later, the corresponding expression was found for the quantum group $\mathrm{U}_{q}(\mathfrak{sl}_{M})$ with arbitrary $M$ [10], and for the general quantum algebra associated with a finite dimensional simple Lie algebra [11, 12]. The explicit expression for the case of the quantum group associated with the simplest affine Lie algebra $\widehat{\mathfrak{sl}_{2}}$ was derived in the paper [13], and for the general affine Lie algebra in the papers by Damiani [14, 15]. Alternative approach was proposed by Khoroshkin and Tolstoy [16]. In fact, the main difference between the papers [14] and [16] lies in the method of constructing the suitable Poincaré–Birkhoff–Witt basis of the quantum algebra. In the paper [14] the Lusztig automorphisms [17] were used, and the authors of [16] used the notion of the $q$-commutator, see also the paper [18].
@@
By generalizing the defining relations of quantum algebra appropriately, one arrives at quantum algebras associated with the Lie superalgebras [19, 20]. The quantum double construction can be directly generalized to the case of such quantum algebras [21]. In the paper [22], it was used to find an explicit form of the universal $R$-matrix for the quantum algebra $\mathrm{U}_{q}(\mathfrak{sl}_{M|N})$. The Khoroshkin–Tolstoy approach for the quantum algebras associated with finite dimensional superalgebras is presented in the paper [23], and associated with Kac–Moody superalgebras with symmetrizable generalized Cartan matrix in the papers [24, 25, 26].
@@
Explicit expressions for the universal $R$-matrix were used to construct the corresponding integrability objects. Here, $R$-operators [27, 13, 28, 29, 30, 31, 32], monodromy operators and $L$-operators were constructed [33, 31, 32, 34, 35, 1]. The corresponding families of functional relations were found [36, 37, 33, 1, 38, 39, 2, 3]. Among other applications, we would like to mention the description of the hidden fermionic structure of the XXZ model [40, 41, 42], and the derivation of equations satisfied by the reduced density operators of the quantum chains related to the quantum group $\mathrm{U}_{q}(\mathfrak{sl}_{M})$ [43, 44].
@@
The above list of applications of the quantum algebraic approach is by no means exhaustive. It should be noted that, apart from the paper [33], it does not contain applications of this approach to quantum integrable systems related to quantum superalgebras. We have not found in the literature examples of derivation of an integrability object for a quantum integrable system related to a quantum superalgebra in the framework of the quantum algebraic approach, see, however, the paper [45], where some hints of the algebraic approach were used for a special case of a superalgebra. Note that the method based on usage of the Lusztig automorphisms is not applicable here, since we do not know their explicit form. Some necessary modifications of the Khoroshkin–Tolstoy approach are presented in the papers [24, 25, 26], but, unfortunately, not all of them. In this paper, using the Khoroshkin–Tolstoy approach, we derive the explicit form of the $R$-operator for the vector representation of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. As expected, up to normalization and arbitrariness of the grading, we obtain the Perk–Schultz $R$-operator [46]. In fact, the main goal of this paper is to investigate the modification of the Khoroshkin–Tolstoy approach which is necessary for its use in constructing integrability objects for systems related to quantum superalgebras.
@@
In section 2 we give general information on the Lie superalgebras $\mathfrak{gl}_{M|N}$ and $\mathfrak{sl}_{M|N}$, describe their root structure, and present the symmetric bilinear form for $\mathfrak{sl}_{M|N}$. The quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ is introduced and its vector representation is defined in section 3. In the beginning of section 4 we discuss the affinization of the Lie superalgebra $\mathfrak{sl}_{M|N}$, and then define the corresponding quantum loop superalgebra. The Khoroshkin–Tolstoy approach to the construction of Cartan–Weyl generators of this superalgebra is considered in section 5. The expression for the universal $R$-matrix is given in section 6. Finally, defining the evaluation module corresponding to the vector representation of $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ is defined in section 7, we find in section 8 the explicit form of the $R$-operator.
@@
Tensor products of quantum algebras are considered to be suitably completed [47].
@@
We fix two positive integers $M$ and $N$ such that $M,N\geq 1$ and $M\neq N$, and denote by $\mathbb{C}_{M|N}$ the superspace formed by $(M+N)$-tuples of complex numbers with the following grading. An element of $\mathbb{C}_{M|N}$ is even if its last $N$ components are zero, and odd if its first $M$ components are zero.  [^3] See appendix A for relevant definitions and notation. For simplicity, we denote the Lie superalgebra Lie $\mathfrak{gl}(\mathbb{C}_{M|N})$ as $\mathfrak{gl}_{M|N}$. Let $v_{i}$, $i=1,\ldots,M+N$, are the elements of the standard basis of $\mathbb{C}_{M|N}$. By definition,
@@
$\mathbb{E}_{ij}v_{k}=v_{i}\delta_{jk}.$
Hence, the matrices of $\mathbb{E}_{ij}$ with respect to the standard basis of $\mathbb{C}_{M|N}$ are the usual matrix units, and we have
@@
It is also clear that
@@
As a Cartan subalgebra $\mathfrak{k}$ of the Lie superalgebra $\mathfrak{gl}_{M|N}$ we use a subalgebra spanned by the elements $K_{i}=\mathbb{E}_{ii}$, $i=1,\ldots,M+N$, which form its basis. Let $(\Xi_{i})_{i=1,\ldots,M+N}$ be the dual basis of $\mathfrak{k}^{*}$. For $X=\sum_{i=1}^{M+N}c_{i}K_{i}\in\mathfrak{k}$ we have
@@
Hence, $\mathbb{E}_{ij}$, $i\neq j$, is the root vector corresponding to the root $\Xi_{i}-\Xi_{j}$ and the root system of $\mathfrak{gl}_{M|N}$ is the set
@@
$\Pi=\{\Xi_{i}-\Xi_{i+1}\mid i=1,\ldots,M+N-1\}.$
The system of positive roots corresponding to $\Pi$ is
@@
Certainly, the corresponding system of negative roots is $\Delta_{-}=-\Delta_{+}$. Denoting the simple roots as
@@
$\alpha_{ij}=\sum_{k=1}^{j-1}\alpha_{k},\qquad 1\leq i<j\leq M+N-1.$
Note that the root $\alpha_{M}$ is odd, all other simple roots are even, see for the definition [48]. Denoting the parity of a root $\gamma\in\Delta$ by $[\gamma]$, we see that
@@
Identifying an element of $\mathfrak{gl}_{M|N}$ with its matrix with respect to the standard basis of $\mathbb{C}_{M|N}$, we define the supertrace of an element $X=(X_{ij}))_{i,j=1}^{M+N-1}$ as
@@
The special linear Lie superalgebra $\mathfrak{sl}_{M|N}$ is a Lie subsuperalgebra of $\mathfrak{gl}_{M|N}$ formed by the matrices with zero supertrace. As a Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{sl}_{M|N}$ we take the subalgebra of $\mathfrak{k}$ which consists of all elements with zero supertrace. The standard basis of $\mathfrak{h}$ is formed by the matrices $H_{i}$, $i=1,\ldots,M+N-1$, defined by the equation
$H_{i}=K_{i}-(-1)^{[i]+[i+1]}K_{i+1}.$
As the positive and negative roots we take the restriction of the roots of $\mathfrak{gl}_{M|N}$ to $\mathfrak{h}$. The numbers
@@
are the entries of the Cartan matrix $A$ of the Lie superalgebra $\mathfrak{sl}_{M|N}$. Putting
@@
and multiplying $A$ from the left by the diagonal matrix
@@
we get the symmetrizied Cartan matrix $B=(B_{ij})_{i,j=1}^{M+N-1}$.
@@
Let $q$ be the exponential of a complex number $\hbar$, such that $q$ is not a root of unity. We define the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ as a unital associative $\mathbb{C}$-algebra generated by the elements [^4] We use capital letters to distinguish between generators of $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ and $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$.
@@
which obey the corresponding defining relations. The $\mathbb{Z}_{2}$-grading of $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ is defined on generators as
@@
Before giving the explicit form of the defining relations, we introduce the notion of the $q$-supercommutator. The abelian group
@@
is called the root lattice of $\mathfrak{gl}_{M|N}$. We denote by $Q_{+}$ and $Q_{-}$ the sublattices
$Q_{+}=\bigoplus_{i=1}^{M+N-1}\mathbb{Z}_{\geq 0}\,\alpha_{i},\qquad Q_{-}=\bigoplus_{i=1}^{M+N-1}\mathbb{Z}_{\leq 0}\,\alpha_{i}.$
Assuming that
@@
we endow $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ with a $Q$-grading.
For any $X\in\mathrm{U}_{q}(\mathfrak{gl}_{M|N})_{\alpha}$ and $Y\in\mathrm{U}_{q}(\mathfrak{gl}_{M|N})_{\beta}$ we define the $q$-supercommutator by the equation
@@
if $\alpha,\beta\in Q_{+}$, by the equation
$[\![X,\,Y]\!]=YX-(-1)^{[X][Y]}q^{(\,\alpha\,|\,\beta\,)}XY=YX-(-1)^{[\alpha][\beta]}q^{(\,\alpha\,|\,\beta\,)}XY$
if $\alpha,\beta\in Q_{-}$, and by the equation
@@
if $\alpha\in Q_{+}$ and $\beta\in Q_{-}$, or $\alpha\in Q_{-}$ and $\beta\in Q_{+}$.
@@
$q^{0}=1,\qquad q^{X_{1}}q^{X_{2}}=q^{X_{1}+X_{2}},$ (3.1)
$[\![E_{i},\,F_{j}]\!]=\delta_{ij}\frac{q_{i}^{H_{i}}-q_{i}^{-H_{i}}}{q_{i}-q_{i}^{-1}},$ (3.3)
$[\![E_{i},\,E_{j}]\!]=0,\qquad[\![F_{i},\,F_{j}]\!]=0,\qquad\mbox{ if }A_{ij}=0,$ (3.4)
$[\![E_{i},\,[\![E_{i},\,E_{i\pm 1}]\!]]\!]=0,\qquad[\![F_{i},\,[\![F_{i},\,F_{i\pm 1}]\!]]\!]=0,\qquad\mbox{ if }A_{ii}\neq 0,$ (3.5)
$[\![[\![[\![E_{M-1},\,E_{M}]\!],\,E_{M+1}]\!],\,E_{M}]\!]=0,\qquad[\![[\![[\![F_{M-1},\,F_{M}]\!],\,F_{M+1}]\!],\,F_{M}]\!]=0,$ (3.6)
@@
$q_{i}-q^{-1}_{i}=d_{i}(q-q^{-1}).$ (3.7)
The general reference on the defining relations of quantum superalgebras is the paper [20].
An element $a$ of $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ is called a root vector corresponding to a root $\gamma$ of $\mathfrak{gl}_{M|N}$ if $a\in\mathrm{U}_{q}(\mathfrak{gl}_{M|N})_{\gamma}$. In particular, $E_{i}$ and $F_{i}$ are root vectors corresponding to the roots $\alpha_{i}$ and $-\alpha_{i}$, respectively. It is possible to construct linearly independent root vectors corresponding to all roots of $\mathfrak{gl}_{M|N}$. To this end, being inspired by M. Jimbo [49], we introduce the elements $E_{ij}$ and $F_{ij}$, $1\leq i<j\leq M+N$, with the help of the relations
$E_{i,\,i+1}=E_{i},\qquad F_{i,\,i+1}=F_{i},$ (3.8)
$E_{i,\,k+1}=[\![E_{ik},\,E_{k,\,k+1}]\!],\qquad F_{i,\,k+1}=[\![F_{ik},\,F_{k,\,k+1}]\!],\qquad k-i>0.$ (3.9)
Explicitly, the last two equations look as
$E_{i,\,k+1}=E_{ik}\,E_{k,\,k+1}-q^{d_{k}}\,E_{k,\,k+1}\,E_{ik},\qquad F_{i,k+1}=F_{k,\,k+1}\,F_{ik}-q^{-d_{k}}F_{ik}\,F_{k,\,k+1}.$
@@
It is clear that the vectors $E_{ij}$ and $F_{ij}$ correspond to the roots $\alpha_{ij}$ and $-\alpha_{ij}$ respectively. These vectors are linearly independent, and together with the elements $q^{X}$, $X\in\mathfrak{k}$ are called Cartan–Weyl generators of $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$. It appears that the ordered monomials constructed from the Cartan–Weyl generators form a Poincaré–Birkhoff–Witt basis of the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$.
In the present paper we work with the so called vector or defining representation $\pi$ of the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ on the superspace $\mathbb{C}_{M|N}$ defined by the relations
@@
This representation is used in section 7 to construct an evaluation representation of the quantum algebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$.

### v3

@@ Abstract.
The central object of the quantum algebraic approach to the study of quantum integrable models is the universal $R$-matrix, which is an element of the completed tensor product of two copies of a quantum algebra. Various integrability objects are constructed by choosing representations for the factors of this tensor product. There are two approaches to constructing explicit expressions for the universal $R$-matrix. One is based on the use of Lusztig automorphisms, and the other is based on the concepts of normal ordering and $q$-commutator. In the case of a quantum superalgebra, we cannot use the first approach, since we do not know an explicit expression for the Lusztig automorphisms. The second approach can be used, although it requires some modifications. In this article, we present the necessary modification of the method and use it to find an $R$-operator for a quantum integrable system related to the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$.
@@ 1. Introduction
The study of quantum integrable models is based on the constructing the corresponding integrability objects and solving the functional relations they satisfy. [^1] For the terminology used for integrability objects, we refer to the papers [1, 2, 3]. Here, the most productive method is the quantum algebraic approach. [^2] Previously, what we call quantum algebra was usually called quantum group. In fact, this object is an associative algebra, which in a sense is a deformation of the universal enveloping algebra of a Lie algebra. Nowadays, the term quantum algebra is more commonly used, and we adhere to this terminology. The central object of this approach is the universal $R$-matrix, which is an element of the completed tensor product of two copies of the quantum algebra. Various integrability objects are constructed by choosing representations for the factors of this tensor product. The consistent application of the quantum algebraic approach was initiated by Bazhanov, Lukyanov and Zamolodchikov [4, 5, 6]. They considered the quantum version of the KdV theory. Subsequently, this method turned out to be efficient for studying other quantum integrable models.
@@
The general notion of a quantum algebra $\mathrm{U}_{q}(\mathfrak{g})$ was proposed by Drinfeld and Jimbo [7, 8] for the case when $\mathfrak{g}$ is a Kac–Moody algebra with a symmetrizable generalized Cartan matrix. In the report [9], Drinfeld described the quantum double construction of the universal $R$-matrix and presented an explicit expression for the quantum algebra $\mathrm{U}_{q}({\mathfrak{sl}_{2}})$. Later, the corresponding expression was found for the quantum algebra $\mathrm{U}_{q}(\mathfrak{sl}_{M})$ with arbitrary $M$ [10], and for a general quantum algebra associated with a finite dimensional simple Lie algebra [11, 12]. The explicit expression for the case of a quantum algebra associated with the simplest affine Lie algebra $\widehat{\mathfrak{sl}_{2}}$ was derived in the paper [13], and for the general affine Lie algebra in the papers by Damiani [14, 15]. Alternative approach was proposed by Khoroshkin and Tolstoy [16]. In fact, the main difference between the papers [14, 15] and [16] lies in the method of constructing the suitable Poincaré–Birkhoff–Witt basis of the quantum algebra. In the paper [14] the Lusztig automorphisms [17] were used, and the authors of [16, 18] used to this end the concepts of normal ordering $q$-commutator.
@@
By generalizing the defining relations of quantum algebra appropriately, one arrives at quantum algebras associated with Lie superalgebras [19, 20]. The quantum double construction can be directly generalized to the case of such quantum algebras [21]. In the paper [22], it was used to find an explicit form of the universal $R$-matrix for the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{sl}_{M|N})$. The Khoroshkin–Tolstoy approach for the quantum superalgebras associated with finite dimensional superalgebras is presented in the paper [23], and for the quantum superalgebras associated with Kac–Moody superalgebras with symmetrizable generalized Cartan matrix, in the papers [24, 25, 26].
@@
Explicit expressions for the universal $R$-matrix were used to construct the corresponding integrability objects. Here, $R$-operators [27, 13, 28, 29, 30, 31, 32], monodromy operators, and $L$-operators were constructed [33, 31, 32, 34, 35, 1]. The corresponding families of functional relations were found [36, 37, 33, 1, 38, 39, 2, 3]. Among other applications of the quantum algebraic approach, we would like to mention the description of the hidden fermionic structure of the XXZ model [40, 41, 42], and the derivation of equations satisfied by the reduced density operator of the quantum chain related to the quantum algebra $\mathrm{U}_{q}(\mathfrak{sl}_{M})$ [43, 44].
@@
The above list of applications of the quantum algebraic approach is by no means exhaustive. It should be noted that, apart from the paper [33], it does not contain applications of this approach to quantum integrable systems related to quantum superalgebras. We have not found in the literature examples of derivation of an integrability object for a quantum integrable system related to a quantum superalgebra in the framework of the quantum algebraic approach, see, however, the paper [45], where some hints of the quantum algebraic approach were used for a special case of a quantum superalgebra. Note that the method based on usage of the Lusztig automorphisms is not applicable here, since we do not know their explicit form. Some necessary modifications of the Khoroshkin–Tolstoy approach are presented in the papers [24, 25, 26], but, unfortunately, not all of them. In this paper, using the Khoroshkin–Tolstoy approach, we derive the explicit form of the $R$-operator for the evaluation vector representation of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. As expected, up to normalization and arbitrariness in the grading, we obtain the Perk–Schultz $R$-operator [46]. In fact, the main goal of this paper is to investigate the modification of the Khoroshkin–Tolstoy approach which is necessary for its use in constructing integrability objects for systems related to quantum superalgebras.
@@
In section 2 we give general information on the Lie superalgebras $\mathfrak{gl}_{M|N}$ and $\mathfrak{sl}_{M|N}$, describe their root structure, and present the symmetric bilinear form for the superalgebra $\mathfrak{sl}_{M|N}$. The quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ and its vector representation are defined in section 3. In the beginning of section 4 we discuss the affinization of the Lie superalgebra $\mathfrak{sl}_{M|N}$, and then define the corresponding quantum superalgebra. The Khoroshkin–Tolstoy approach to the construction of Cartan–Weyl generators of this superalgebra is considered in section 5. The expression for the universal $R$-matrix is given in section 7. Finally, the evaluation module corresponding to the vector representation of $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ is defined in section 6, and the explicit form of the corresponding $R$-operator is found in section 8.
@@
Tensor products of quantum algebras are considered to be suitably completed, see the book [47].
@@
We fix two positive integers $M$ and $N$ such that $M,N\geq 1$ and $M\neq N$, and denote by $\mathbb{C}_{M|N}$ the superspace [^3] See appendix A for relevant definitions and notation. formed by $(M+N)$-tuples of complex numbers with the following grading. An element of $\mathbb{C}_{M|N}$ is even if its last $N$ components are zero, and odd if its first $M$ components are zero. For simplicity, we denote the Lie superalgebra $\mathfrak{gl}(\mathbb{C}_{M|N})$ as $\mathfrak{gl}_{M|N}$. We denote by $v_{i}$, $i=1,\ldots,M+N$, the elements of the standard basis of $\mathbb{C}_{M|N}$. By definition,
@@
$\mathbb{E}_{ij}v_{k}=v_{i}\delta_{jk},$
form a basis of the Lie superalgebra $\mathfrak{gl}_{M|N}$. It is clear that the matrices of $\mathbb{E}_{ij}$ with respect to the standard basis of $\mathbb{C}_{M|N}$ are the usual matrix units, and we have
@@
It is also evident that
@@
As the Cartan subalgebra $\mathfrak{k}$ of the Lie superalgebra $\mathfrak{gl}_{M|N}$ we take the subalgebra spanned by the elements $K_{i}=\mathbb{E}_{ii}$, $i=1,\ldots,M+N$, which form its basis. Denote by $(\Xi_{i})_{i=1,\ldots,M+N}$ the dual basis of the space $\mathfrak{k}^{*}$. For $X=\sum_{i=1}^{M+N}c_{i}K_{i}\in\mathfrak{k}$ we have
@@
Hence, $\mathbb{E}_{ij}$, $i\neq j$, is a root vector corresponding to the root $\Xi_{i}-\Xi_{j}$ and the root system of $\mathfrak{gl}_{M|N}$ is the set
@@
$\Pi=\{\Xi_{i}-\Xi_{i+1}\mid i=1,\ldots,M+N-1\},$
then the system of positive roots corresponding to $\Pi$ is
@@
Certainly, the corresponding system of negative roots is $\Delta_{-}=-\Delta_{+}$. Denoting
@@
$\alpha_{ij}=\sum_{k=1}^{j-1}\alpha_{k},\qquad 1\leq i<j\leq M+N.$
Note that the root $\alpha_{M}$ is odd, all other simple roots are even, see for the definition the paper [48]. Denoting the parity of a root $\gamma\in\Delta$ by $[\gamma]$, we see that
@@
Like any root system, $\Delta$ defines a strict partial order $\prec$ on $\mathfrak{h}^{*}$. Here, for $\alpha,\beta\in\mathfrak{h}^{*}$ we have $\beta\prec\alpha$ if and only if $\alpha-\beta$ is the sum of positive roots.
Identifying an element of the Lie superalgebra $\mathfrak{gl}_{M|N}$ with its matrix with respect to the standard basis of the superspace $\mathbb{C}_{M|N}$, we define the supertrace of an element $X=(X_{ij})_{i,j=1}^{M+N}$ by the relation
@@
The Lie superalgebra $\mathfrak{sl}_{M|N}$ is a Lie subsuperalgebra of the Lie superalgebra $\mathfrak{gl}_{M|N}$ formed by the elements with zero supertrace. As the Cartan subalgebra $\mathfrak{h}$ of the Lie superalgebra $\mathfrak{sl}_{M|N}$ we take the Lie subsuperalgebra $\mathfrak{h}$ of the Lie superalgebra $\mathfrak{k}$, which consists of the elements with zero supertrace. The elements $h_{i}\in\mathfrak{h}$, $i=1,\ldots,M+N-1$, defined by the equation
$h_{i}=K_{i}-(-1)^{[i]+[i+1]}K_{i+1}$
form a basis of $\mathfrak{h}$. As the positive and negative roots we take the restriction of the positive and negative roots of the Lie superalgebra $\mathfrak{gl}_{M|N}$ to $\mathfrak{h}$. The numbers
@@
are the entries of the Cartan matrix $A=(a_{ij})_{i,j=1}^{M+N-1}$ of the Lie superalgebra $\mathfrak{sl}_{M|N}$. The explicit form of the Cartan matrix $A$ is shown in figure 1.
$A=\left[\begin{array}[]{rrrrr|r|rrrrr}2&-1&&&&\\ -1&2&-1&&&\\ &\ddots&\ddots&\ddots&&\\ &&\ddots&\ddots&\ddots&\\ &&&-1&2&-1\\ \hline\cr&&&&-1&0&1\\ \hline\cr&&&&&-1&2&-1\\ &&&&&&\ddots&\ddots&\ddots\\ &&&&&&&\ddots&\ddots&\ddots\\ &&&&&&&&-1&2&-1\\ &&&&&&&&&-1&2\\ \end{array}\right]$
Figure 1. The Cartan matrix of the Lie superalgebra $\mathfrak{sl}_{M|N}$ with distinguished $M$th row and $M$th column.
Putting
@@
and multiplying the matrix $A$ from the left by the diagonal matrix
@@
we get the symmetrizied Cartan matrix $B=DA=(b_{ij})_{i,j=1}^{M+N-1}$. For the explicit form of the matrix $B$ see figure 2.
$B=\left[\begin{array}[]{rrrrr|r|rrrrr}2&-1&&&&\\ -1&2&-1&&&\\ &\ddots&\ddots&\ddots&&\\ &&\ddots&\ddots&\ddots&\\ &&&-1&2&-1\\ \hline\cr&&&&-1&0&1\\ \hline\cr&&&&&1&-2&1\\ &&&&&&\ddots&\ddots&\ddots\\ &&&&&&&\ddots&\ddots&\ddots\\ &&&&&&&&1&-2&1\\ &&&&&&&&&1&-2\\ \end{array}\right]$
Figure 2. The symmetrized Cartan matrix of the Lie superalgebra $\mathfrak{sl}_{M|N}$ with distinguished $M$th row and $M$th column.
@@
We define the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ as a unital associative $\mathbb{C}$-superalgebra generated by the elements [^4] We use capital letters to distinguish between generators of the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ and the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$.
@@
which obey the corresponding defining relations. The $\mathbb{Z}_{2}$-grading of the superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ is defined on generators as
@@
Before giving the explicit form of the defining relations, introduce the notion of the $q$-supercommutator. The abelian group
@@
is called the root lattice of the Lie superalgebra $\mathfrak{gl}_{M|N}$. Assuming that
@@
we endow $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ with a $Q$-grading. Now, for any elements $X\in\mathrm{U}_{q}(\mathfrak{gl}_{M|N})_{\alpha}$ and $Y\in\mathrm{U}_{q}(\mathfrak{gl}_{M|N})_{\beta}$ define the $q$-supercommutator by the equation
@@
if $\alpha,\beta\succ 0$, by the equation
$[\![X,\,Y]\!]=XY-(-1)^{[X][Y]}q^{(\,\alpha\,|\,\beta\,)}YX=XY-(-1)^{[\alpha][\beta]}q^{(\,\alpha\,|\,\beta\,)}XY$
if $\alpha,\beta\prec 0$, and by the equation
@@
if $\alpha\prec 0$ and $\beta\succ 0$, or $\alpha\succ 0$ and $\beta\prec 0$.
@@
$q^{0}=1,\qquad q^{X_{1}}q^{X_{2}}=q^{X_{1}+X_{2}},$
$q^{X}E_{i}q^{-X}=q^{\langle\alpha_{i},\,X\rangle}E_{i},\qquad q^{X}F_{i}q^{-X}=q^{-\langle\alpha_{i},\,X\rangle}F_{i},$
$[\![E_{i},\,F_{j}]\!]=\delta_{ij}\frac{q_{i}^{H_{i}}-q_{i}^{-H_{i}}}{q_{i}-q_{i}^{-1}},$
$[\![E_{i},\,E_{j}]\!]=0,\qquad[\![F_{j},\,F_{i}]\!]=0,\qquad\mbox{ if }(\,\alpha_{j}\,|\,\alpha_{i}\,)=0,$
$[\![E_{i},\,[\![E_{i},\,E_{i-1}]\!]]\!]=0,\qquad[\![[\![F_{i-1},\,F_{i}]\!],\,F_{i}]\!]=0,\qquad\mbox{ if }i\neq 1\mbox{ and }(\,\alpha_{i}\,|\,\alpha_{i}\,)\neq 0,\hskip 28.00006pt$
$[\![E_{i},\,[\![E_{i},\,E_{i+1}]\!]]\!]=0,\qquad[\![[\![F_{i+1},\,F_{i}]\!],\,F_{i}]\!]=0,\qquad\mbox{ if }i\neq M+N\mbox{ and }(\,\alpha_{i}\,|\,\alpha_{j}\,)\neq 0,$
$[\![[\![[\![E_{M-1},\,E_{M}]\!],\,E_{M+1}]\!],\,E_{M}]\!]=0,\qquad[\![F_{M},\,[\![F_{M+1},\,[\![F_{M},\,F_{M-1}]\!]]\!]]\!]=0,$
@@
$q_{i}-q^{-1}_{i}=d_{i}(q-q^{-1}).$
A general description of the defining relations of quantum superalgebras is given in the paper [20].
An element $a\in\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ is called a root vector corresponding to a root $\gamma$ of the Lie superalgebra $\mathfrak{gl}_{M|N}$ if $a\in\mathrm{U}_{q}(\mathfrak{gl}_{M|N})_{\gamma}$. In particular, $E_{i}$ and $F_{i}$ are root vectors corresponding to the roots $\alpha_{i}$ and $-\alpha_{i}$, respectively. It is possible to construct linearly independent root vectors corresponding to all roots of the Lie superalgebra $\mathfrak{gl}_{M|N}$. To this end, following by Jimbo [49], we introduce the elements $E_{ij}$ and $F_{ij}$, $1\leq i<j\leq M+N$, with the help of the relations
$E_{i,\,i+1}=E_{i},\qquad F_{i,\,i+1}=F_{i},$
$E_{i,\,k+1}=[\![E_{ik},\,E_{k,\,k+1}]\!],\qquad F_{i,\,k+1}=[\![F_{k,\,k+1},\,F_{ik}]\!],\qquad k>i.$
In a more explicit form, the last two equations look as
$E_{i,\,k+1}=E_{ik}\,E_{k,\,k+1}-q_{k}\,E_{k,\,k+1}\,E_{ik},\qquad F_{i,k+1}=F_{k,\,k+1}\,F_{ik}-q_{k}^{-1}F_{ik}\,F_{k,\,k+1}.$
@@
It is clear that $E_{ij}$ and $F_{ij}$ are root vectors corresponding to the roots $\alpha_{ij}$ and $-\alpha_{ij}$, respectively. These vectors are linearly independent, and together with the elements $q^{X}$, $X\in\mathfrak{k}$, are called the Cartan–Weyl generators of the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$. It appears that the ordered monomials constructed from the Cartan–Weyl generators form a Poincaré–Birkhoff–Witt basis of the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$.
In the present paper we will work with the so called vector representation $\pi$ of the quantum superalgebra $\mathrm{U}_{q}(\mathfrak{gl}_{M|N})$ on the superspace $\mathbb{C}_{M|N}$ defined by the relations
@@
This representation will be used in section 6 to construct the evaluation vector representation of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$.